--++ vidyo

vidyo link PIN 1055

Action Items:


Tom : Should we maintain minutes ? Yes.

Paolo : Timing Layer Days 22nd and 23rd March. Massimo is going there for one day. The detector was approved by the LHCC and research board. TDR due in ~ one year. Anna : workshop at CERN on slice test for DT upgrade on Wed. 14th (wukl be held during long shutdown). One week test beam at the gif++ on irradiated MB2 chamber (25 april , may 2). Experienced shifters are needed. Paolo : EPR ? Yes.

Framework & production


B physics

Martino : Alessio will presente the L1 trigger paths strategy at the DPG today. Alberto : soft muon ID implementaion on course. My comments from the responsible, should coneverge quite soon. Franco : will we have something by the brussels internal meeting ? Yes (A).

Charged Higgs to top-beauty

No news.

Dibosons, VZ resonance (V=W,Z)

Lisa' analysis (2q2nu) submitted to JHEP. Final reading for 2l2q on Thursday, working in order to prepared a full compilation of searches.

4B full hadronic

Move to standard model BDT, last step before preapproval. Limits got worse, we are within 2 sigmas.

Heavy New Fermions

LHE production for Heavy Neutrino at 14TeV approved by EXO-MC convener. LHE being processed with Delphes+Pythia by UPSG (0PU already available).

Higgs Dark Matter

Siew Yan : sensitivity in 2016 , 2017: need to perfrom analysis also on 2017. SY works on documentation and a proxy analysis on the monojet, then (by fall) will move away keeping just a managerial role. Move to higgssthrhahllung.

Same sign dileptons


Roberto is on shift (even on Eater !) Marina : explore the possibily to be FOG coordination from remote.


Jacopo : worked with Andrea Triossi. Firmware assembled, but the Spanish guys asked for a different setup. Wrapping the unpacker is a bit delayed, by end this week the two parts should start to be integrated.


DT : ongoing preparation for pp collisions. MB1 chambers on the external wheel will run at lowr voltage (3550, 3500 ) to drive less current with unaffected efficiency.

Also sheild on top MB4 chambers of all wheels.

Luminosity & Bril

Timing Layer


Paolo : what's the status with the inner layer problem (DC/DC converter issue) ? Not solved as far as I know (Jacopo). Will be discussed during the CMS week.


-- FrancoSimonetto - 2018-03-12

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2018-03-12 - MatteoPresilla
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback